Will some one please help me with this word problem, I answered a) but am not even sure if i have that right, wil someone help me please!
1. You have recently found a location for your bakery and have begun implementing the first phases of your business plan. Your budget consists of an $80,000 loan from your family and a $38,250 small business loan. These loans must be repaid in full within 10 years.
a) What integer would represent your total budget? $118,250
b) Twenty-five percent of your budget will be used to rent business space and pay for utilities. Write an algebraic expression that indicates how much money will be spent on business space and utilities. Do not solve.
c) How much money will rent and utilities cost? Explain how you arrived at this answer.
d) Suppose an investor has increased your budget by $22,250. The investor does not need to be repaid. Rather, he becomes part owner of your business.
Will the investor contribute enough money to meet the cost of rent and utilities? Support your answer, and write an equation or inequality that illustrates your answer.
e) This equation illustrates your remaining funds after paying for rent and utilities. How much money is left? Explain how you arrived at your answer.
$38,250 + $80,000+ $22,250-0.25($80,000 + $38,250) =
¬

Answer by vleith(2983)

a is correct
b) rentandUtiliets = Budget*0.25
c) rentandUtiliets = Budget*0.25
rentAndUtilities = 118,250 * 0.25 = 29,562.50
d)No, the investor added 22,250 and the rent is 29,562.50
22,250 < Budget*0.25
e) money left is 0.75(118,250) + 22,250 = 110,937.50
The money left is the initial set of loans money less the rent. That is (1-0.25)* 118250. Then we have to add in the partners investment of 22,250. That leaves 110.937.50